OpenStudy (anonymous):

Find the product of the complex numbers. Express your answer in trigonometric form. z1=3(cos(pi/4)+i sin(pi/4)) z2=7(cos(3pi/8)+i sin(3pi/8))

OpenStudy (tkhunny):

Have you considered multiplying the leading coefficients and adding the angles?

OpenStudy (anonymous):

so 3*7?

OpenStudy (tkhunny):

Now, the second part.

OpenStudy (anonymous):

multiply pi/4*pi/4?

OpenStudy (tkhunny):

Read my first post.

OpenStudy (anonymous):

Add pi/4+pi/4?

OpenStudy (tkhunny):

Why would you do that? When you came up with 3*7, you took a value from each of z1 and z2. You should do the same with the addition.

OpenStudy (anonymous):

you add 3 and 7? is that what you're saying?

OpenStudy (tkhunny):

No, Please don't reverse correct things you did already. Multiply the leading constants: 3*7 -- Done. Don't change that. Add the angles: ?? + ??

OpenStudy (anonymous):

ok I think I understand I use the unit circle to figure out the angle then I add?

OpenStudy (tkhunny):

?? The angles are given. Why do you need a unit circle? \(\dfrac{\pi}{4}+\dfrac{3\pi}{8}\) -- Go!

OpenStudy (anonymous):

5pi/8

OpenStudy (tkhunny):

Now, if this happens to be greater than \(2\pi\), you might need that Unit Circle. What is your final answer? You have all the pieces.

OpenStudy (anonymous):

So do I do 21cos(5pi/8)+i sin(5pi/8) ?

OpenStudy (tkhunny):

Not quite. The parentheses are not optional.

OpenStudy (anonymous):

21(cos(5pi/8)+i sin(5pi/8))

OpenStudy (tkhunny):

Perfect.

OpenStudy (anonymous):

okay let me solve

OpenStudy (tkhunny):

Solve what? You are done. Move on to the next one.

OpenStudy (anonymous):

oh that's it?

OpenStudy (tkhunny):

That's all there is to it. :-)

OpenStudy (anonymous):

thanks